OpenStudy (anonymous):

9^(x+1)=27^(x-1) Can someone PLZ explain how to solve this?

OpenStudy (turingtest):

write 9 and 27 as powers of 3

OpenStudy (turingtest):

then use the rule\[(x^a)^b=x^{ab}\]

OpenStudy (turingtest):

9=3^? 27=3^?

OpenStudy (turingtest):

please talk to me so I can help you...

OpenStudy (anonymous):

i'm still not understanding. What do u mean write 9 & 27 as powers of 3?

OpenStudy (turingtest):

3 to what power equals 9 ?

OpenStudy (anonymous):

3^2=9

OpenStudy (turingtest):

yes, good :) and 27=3^?

OpenStudy (anonymous):

3^3

OpenStudy (turingtest):

good, so we have\[\large 9^{x+1}=27^{x-1}\]\[\large(3^2)^{x+1}=(3^3)^{x+1}\]and now we use the rule\[\large(x^a)^b=x^{ab}\]

OpenStudy (anonymous):

ok now how do i use that rule?

OpenStudy (turingtest):

in your case we apply the rule on the left as\[x=3,~a=2,~b=x+1\]so we get\[\large(x^a)^b=x^{ab}\]\[\large(3^2)^{x+1}=3^{2(x+1)}=3^{2x+2}\]try to meditate on that until it makes sense...

OpenStudy (turingtest):

on the other side we do the same thing with\[\large(3^3)^{x-1}=?\]do this one yourself based on the way I did the other side above

OpenStudy (anonymous):

I got \[3^{3x+4}\] is that right?

OpenStudy (turingtest):

not quite\[\large(3^3)^{x-1}=3^{3(x-1)}=?\]

OpenStudy (anonymous):

\[3^{3x-3}\]?

OpenStudy (turingtest):

yes, very good :)

OpenStudy (turingtest):

so now look at what we have on both sides...

OpenStudy (waleed_imtiaz):

U should make the bases same in order to make the exponents same and then solve for x...... write 27 as 3^3(x-1) and 9 as 3^2(x+1) now u can see..... 3 is the base on both the sides and if the bases are the same, then the exponents must be equal..... so...... 2(x+1)=3(x-1).... solve for x now.....

OpenStudy (turingtest):

\[\large 9^{x+1}=27^{x-1}\]\[\large(3^2)^{x+1}=(3^3)^{x-1}\]\[3^{2x+2}=3^{3x-3}\]now you can take the logarithm of both sides, tight?

OpenStudy (anonymous):

\[3^{2x+2}=3^{3x-3}\] Then i did 2x+2=3x-3 & got x=-6. Is that right?

hero (hero):

I was just about to say...there's no need to take logs of both sides.

OpenStudy (turingtest):

2x+2=3x-3 ^^correct, but you did not solve this right

OpenStudy (turingtest):

@Hero I suppose you mean to say that you can just say "the exponents have to be equal" and look at it that way?

hero (hero):

yeah @TuringTest, I explained that to her earlier that's why she knew to set the exponents equal, but her value of x is incorrect.

OpenStudy (turingtest):

right, gotchya

OpenStudy (anonymous):

one sec et me try again

OpenStudy (turingtest):

so @joyce153 try again to solve 2x+2=3x-3

hero (hero):

I never said that her value of x was correct. I simply said there was no need to take logs

OpenStudy (turingtest):

I know, I got it, thanks. I only made the point about taking logs because I assumed she was in a section in which one is learning to use logarithms

OpenStudy (anonymous):

is x=-5?

OpenStudy (turingtest):

closer but still a bit off careful with your signs

hero (hero):

How did you get the negative?

hero (hero):

You should probably show all your steps

OpenStudy (anonymous):

ok i did 2x+2=3x-3 then i did 2x-3x+2=3x-3x-3=-x+2=-3 -x+2-2=-3-2=-x=-5

OpenStudy (turingtest):

oh nevermind

OpenStudy (turingtest):

it just confuses when you write it that way but you have everything right anyway -x=-5 so x=?

hero (hero):

You want to simplify in such a way where the value of x ends up being positive. If it is negative, then you're not done solving for x.

OpenStudy (anonymous):

a neg & a neg = a pos. right? so it be x=5?

OpenStudy (turingtest):

yes

hero (hero):

Here's how I usually simplify these @joyce153, 2x+2=3x-3 I want x term to be positve so I subtract the smaller term, 2x from both sides, then add 3 to both sides to get 3 + 2 = 3x - 2x 5 = x That avoids having to deal with negatives at the end.

OpenStudy (anonymous):

:( i can only give one of u a meddle but THX @Hero & @TuringTest !!

OpenStudy (turingtest):

indeed, and you should write your work thee way Hero has shown above much easier to follow for both you and the person grading your work

OpenStudy (anonymous):

ok!